OpenStudy (anonymous):

Solve x + 2y = 134x + 8y = -9 (3, 5) (-1, 7) No Solutions Infinitely Many Solutions (Basically I've been learning the wrong material and do not know how to go about this material.)

OpenStudy (anonymous):

Please stop writing these as compound equalities when they are not.

OpenStudy (anonymous):

Now how are you gonna solve them...is it 134 or 13 and 4

OpenStudy (anonymous):

No answer either way

OpenStudy (anonymous):

That is how they are given to me, @CliffSedge

OpenStudy (anonymous):

Are there many solutions or none at all?

OpenStudy (anonymous):

134x*

OpenStudy (anonymous):

Here is a multiple choice question for you: What is your system of equations? A) x+2y=1, 34x+8y=-9 B) x+2y=13, 4x+8y=-9 C) x+2y=134, x+8y=-9 D) None of the above

OpenStudy (anonymous):

None of these types has a right answer in the choices given in the question!!!

OpenStudy (anonymous):

Where are you seeing these questions? If they are from an online problem set, you should alert the web site developer that they are getting output errors on their pages.

OpenStudy (anonymous):

here is my guess solve \[ x + 2y = 13\]\[4x + 8y = -9\] lines are parallel so there is no solution

OpenStudy (anonymous):

Well, may be you are right @satellite73

OpenStudy (anonymous):

if it really says \[ x + 2y = 134x + 8y = -9\] then there is a massive typo

OpenStudy (anonymous):

We've seen these same kinds of typos several other times, Mr. satellite.

OpenStudy (anonymous):

It is a systematic flaw in wherever @InsanelyChaotic is drawing these questions from.

OpenStudy (anonymous):

That's how they are brought up to me in context

OpenStudy (anonymous):

are they from the notes given to you or a text book or a website?

OpenStudy (anonymous):

Again, where are you seeing these, @InsanelyChaotic ? If they are from an online course, you need to send an email to whoever is in charge so they can go in and fix these errors.

OpenStudy (anonymous):

The answer I had was "no solutions" all along. I just needed to make sure that there werent a various amount of solutions

OpenStudy (anonymous):

Yes, an online course

OpenStudy (anonymous):

It's from a website

OpenStudy (anonymous):

Do you understand from satellite's astute guess that the system is x+2y=13, 4x+8y=-9, why there is no solution?

OpenStudy (anonymous):

All of these have errors in it, how is the professor to say I have to solve such problems?

OpenStudy (anonymous):

Ahhhh..then do email the admin of that website

OpenStudy (anonymous):

Yes, when he explained it, it made sense. I needed background info as to how it wasn't a solution and received just that. thanks @satellite73

OpenStudy (anonymous):

tell your professor that the online course shows massive typos...

OpenStudy (anonymous):

Call or email your professor and tell that person what is going on and have it fixed. If other students are having this same issue, it can cause a major problem for people's grades.

OpenStudy (anonymous):

This one is no different: Solve 2x + 5y = -133x - 4y = -8 Answer (4, 1) (-4, 1) (4, -1) (-4, -1)

OpenStudy (anonymous):

Ok, I'm about to email, but what exactly do I say?

OpenStudy (anonymous):

Say that the systems of equations are getting smushed together so you can't tell where one ends and the other begins.

OpenStudy (anonymous):

Copy an example to show as well.

OpenStudy (anonymous):

its 2x+5y= -13 and 3x-4y= -8

OpenStudy (anonymous):

Is that right, though? To say such thing? The professor could easily say that it's a part of math and that part of solving is understanding where it begins and ends

OpenStudy (anonymous):

Which specific example do I copy? The one that satellite offered?

OpenStudy (anonymous):

No he is wrong if he says so...You are not to find out the starting and ending points of the equation from the answer choices provided...

OpenStudy (anonymous):

For 2x + 5y = -133x - 4y = -8, you have to guess where the break is supposed to be: 2x + 5y = -1 | 33x - 4y = -8 2x + 5y = -13 | 3x - 4y = -8 2x + 5y = -133 | x - 4y = -8 In the meantime before it gets fixed, you'll have to solve all of the above to find the right answer. I'm guessing that it's the one in the middle since so far all your coefficients have been single-digit numbers.

OpenStudy (anonymous):

Yep that's what i said and you are right..

OpenStudy (anonymous):

"Is that right, though? To say such thing? The professor could easily say that it's a part of math and that part of solving is understanding where it begins and ends" Yes, it is right to say such a thing. It is the only ethical thing to do. If the professor said such a silly thing in response, then he should be fired.

OpenStudy (anonymous):

I'm emailing as we speak, thank you guys

OpenStudy (anonymous):

go with the example which we are discussing now...but its okay to send any example

OpenStudy (anonymous):

I wish I could hug all of you haha

OpenStudy (anonymous):

Alright, definitely

OpenStudy (anonymous):

You can provide all of the examples that you've seen so far. You can say, "Here in problem #__ the system of equations looked like ________, and it was very confusing. Same thing here in problem #___. It looked like _________."

OpenStudy (anonymous):

*HUGS* @InsanelyChaotic You've been a real trooper dealing with this madness. It's fun for me 'cause I'm a huge nerd, but you shouldn't have to deal with such confusion.

OpenStudy (anonymous):

@CliffSedge I agree once again!!!!

OpenStudy (anonymous):

Thank you very much @CliffSedge And people wonder why I hate math.....

OpenStudy (anonymous):

So, are you up with any such UNIQUE questions left to deal with?!

OpenStudy (anonymous):

Yes, far too many

OpenStudy (anonymous):

Here is the email, keep in mind that talking about math, much less pointing out math errors is not my strong suit:

OpenStudy (anonymous):

I'm running in some trouble pertaining to the second module. It's taking me longer because I'm getting mixed up after taking notes on each lesson. Also, I've been hitting the "save" button more than I have been pressing the "submit" button, but I am not far behind. I second-guess a lot of my work. Anyway, what I came to you for was that when I'm going thru some of the quizzes I'm seeing things that don't seem accurate. Bear with me. Solve x + 2y = 134x + 8y = -9 (3, 5) (-1, 7) No Solutions Infinitely Many Solutions (This is seen as a compound equality. My guess: x + 2y = 13 4x + 8y = -9) I see a typo if it says x + 2y = 134x + 8y = -9

OpenStudy (anonymous):

Just try to identify the right equations or if I were you, I would leave this bloody problems aside and continue on some other work!!

OpenStudy (anonymous):

But I want to proof the professor wrong!!!! ;)

OpenStudy (anonymous):

Had the point but you could have mentioned the number of such problems you dealt with.

OpenStudy (anonymous):

Now....you can't do that unless he tries to justify the equations represented in the module even if they are wrong

OpenStudy (anonymous):

I need specific details as to why it's wrong to put in the email

OpenStudy (anonymous):

"And people wonder why I hate math....." Don't hate the math, hate the poorly-stated problems.

OpenStudy (anonymous):

You haven't met this man, he'll do and say anything to proof he's right

OpenStudy (anonymous):

@CliffSedge Fair enough, but it's not my strong suit

OpenStudy (anonymous):

I don't think he can do that if we know what is right

OpenStudy (anonymous):

If he gives you a hard time about it, give me his email address and I'll lay the mathematical smack down on him.

OpenStudy (anonymous):

Right-OOO @CliffSedge

OpenStudy (anonymous):

What would you say in the mathematical smack-down? hahaha

OpenStudy (anonymous):

Specific examples.....? Or is the email good?

OpenStudy (anonymous):

I'd outline any and all of his rebuttals and go through them one-by-one, showing with mathematical proof why they are wrong.

OpenStudy (anonymous):

the email is good..Don't worry about that!!!

OpenStudy (anonymous):

I think your email, as is, is fine. Make sure to mention that all of the systems-of-equations questions that you've seen have had the same flaw.

OpenStudy (anonymous):

English is my strong suit, I worry.....! ;)

OpenStudy (anonymous):

Shall I tell you all what the response is when I receive it?

OpenStudy (anonymous):

hahaha..the prose is tight..you did it well

OpenStudy (anonymous):

Oh, then you'll probably want to fix a few grammatical errors: "I'm running in some trouble ..." Should be "I'm running in to some trouble." "...when I'm going thru some of the quizzes..." misspelling of 'through.'

OpenStudy (anonymous):

My error, I typed it in frustration and in an overall rush

OpenStudy (anonymous):

"Thru" is the lazy way to spell "through"

OpenStudy (anonymous):

Well, I knew you were not to make that error if you had enough time and calm mind while sending that email..

OpenStudy (anonymous):

Sarcasm? haha

OpenStudy (anonymous):

This has me so worked up

OpenStudy (anonymous):

hahaha...!!

OpenStudy (anonymous):

okay..its an absolute fun talking and working with you both PARTNERS...its 1am here and i've to go to bed..GOOD NIGHT

OpenStudy (anonymous):

I think a cold drink and a walk outside is the correct next step in this process.

OpenStudy (anonymous):

Thank you @AbhimanyuPudi !!!!!!

OpenStudy (anonymous):

(or a cold drink and a lie down in a bed in Abhimanyu's case)

OpenStudy (anonymous):

If only I could lie down.... more math...!

OpenStudy (anonymous):

my mom hits me with my cricket bet if i try to have a cold drink now!!!

OpenStudy (anonymous):

*bat

OpenStudy (anonymous):

hahahahahahahhaha

OpenStudy (anonymous):

your mom sounds insane

OpenStudy (anonymous):

No no...i was just kidding..my mom is asleep and she doesn't know anything of what i am doing now...

OpenStudy (anonymous):

Get some sleep :) you earned it, thank you again

OpenStudy (anonymous):

you are welcome...bye